OpenStudy (anonymous):

solve for x and y x-2y=3 .......eqn 1 -2x+y=3 .....eqn 2

OpenStudy (anonymous):

use elimination or substitution method..your value of y is -3

OpenStudy (anonymous):

your value of x is -3 also....

OpenStudy (anonymous):

by the way either of the two method you use, you get the same asnwer, to check if it is right, try substituting to the original equation...

OpenStudy (jmark):

"By making use of elimination method we solve the equation to find the value of x and y . x-2y=3 ......(1) -2x+y=3 .....(2) multiply equation(1) by 2 2x-4y=6 ......(1) -2x+y=3 .....(2) ▬▬▬▬▬▬ -3y = 9 ▬▬▬▬▬▬ y = -3 assign y value in equation (1) x-2(-3)=3 x +6 = 3 x = 3-6 x =- 3 the value of x = -3 and y = -3 ICSE board syllabus for class 8 - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-8.html